### v1

# Efficient Model Editing with Task Vector Bases: A Theoretical Framework and Scalable Approach

Abstract
Task vectors, which are derived from the difference between pre-trained and fine-tuned model weights, enable flexible task adaptation and model merging through arithmetic operations such as addition and negation. However, existing approaches often rely on heuristics with limited theoretical support, often leading to performance gaps comparing to direct task fine tuning. Meanwhile, although it is easy to manipulate saved task vectors with arithmetic for different purposes, such compositional flexibility demands high memory usage, especially when dealing with a huge number of tasks, limiting scalability. This work addresses these issues with a theoretically grounded framework that explains task vector arithmetic and introduces the task vector bases framework. Building upon existing task arithmetic literature, our method significantly reduces the memory cost for downstream arithmetic with little effort, while achieving competitive performance and maintaining compositional advantage, providing a practical solution for large-scale task arithmetic.
Machine Learning, ICML
1 Introduction
Figure 1: Saving one task vector for each task can lead to high memory cost and suboptimal merging performance as the number of tasks increases. Instead, we can construct the affinity graph of task vectors for clustering, identify the clusters, and then apply existing merging methods on them to create task vector bases.
Task vector (Ilharco et al., 2022) is a practical model editing and merging technique that navigates the weight space of pretrained models. These vectors provide a direction that enhances task-specific knowledge in parameter space by subtracting the pretrained model weights from the updated parameters after fine-tuning a specific task. Beyond their simplicity, task vectors have an intriguing property: they can be manipulated through arithmetic operations such as addition and negation, allowing for the composition of models with tailored behaviors or hybrid task capabilities.
However, existing approaches mainly rely on heuristics and lack theoretical justifications (Yang et al., 2024), leaving key questions about their principles and limitations unanswered. The foundations of model merging techniques are nascent. Some existing analyses make impractical assumptions, such as disjoint task support (Ortiz-Jimenez et al., 2024; Xiong et al., 2024) or convexity (Zhou et al., 2024b; Tao et al., 2024), which fail to provide a convincing explanation of the underlying mechanisms of task arithmetic.
Furthermore, the scalability of task vector-based approaches for model editing faces several challenges. One major issue is that merging models trained on multiple tasks often results in a performance drop compared to training models independently or using multi-task joint training, especially as the number of tasks increases (Ilharco et al., 2022; Yang et al., 2024). Additionally, existing model merging methods are inefficient, often requiring substantial memory to store task vectors, which are the same size as the pretrained model. For example, merging 72 fine-tuned ViT-B/32 (Dosovitskiy, 2020) models can require over 200GB of memory (Li et al., 2024; Yang et al., 2024). There were attempts that address this problem either through model sparsification (He et al., 2024; Wang et al., 2024a), alternative optimization algorithms (Li et al., 2024), or assuming all fine-tuned weights lie in a thin Gaussian shell when training task is the same (Jang et al., 2025). However, it is unclear whether the same statement applies to task arithmetic where models are finetuned from diverse tasks. This limits the applicability of such techniques, particularly when handling large-scale models or resource-constrained environments.
In this work, we address these limitations with the following contributions. First, we provide an in-depth theoretical analysis to explain when and why task arithmetic is successful. We establish the connections between key assumptions on task vector norms, task vector similarity, local smoothness, and the generalization error bounds of different task arithmetic operations. We additionally validate our assumptions and theorems with empirical evidence from previous literature and our new setups. Second, as shown in Figure 1, we leverage the similarity between different task vectors and propose a two-step framework to learn a task space with bases: first clustering the task vectors with their similarity matrix, and then merging clustered task vectors as bases with any model merging methods. Through extensive experiments, we show how to do bases arithmetic, including addition, negation and out-of-distribution task generalization with a significant reduction of memory requirement. Our task vector bases approach enables scalable and efficient model editing while preserving strong performance across multiple tasks by fully leveraging the advantage of existing task arithmetic methods in the past literature.
2 Related Work
Theory of Task Arithmetic
Task arithmetic was believed to be related to the hypothesis that fine-tuning over-parameterized models behaves like a neural tangent kernel (Jacot et al., 2018), but this hypothesis contradicts the experimental results in Ortiz-Jimenez et al. (2024) where there exists a nontrivial gap between linear and standard non-linear fine-tuning generalization performance. Therefore, Ortiz-Jimenez et al. (2024) proposed a formal definition of the task arithmetic property, which has been widely adopted by subsequent work (Xiong et al., 2024) for analysis, yet under the assumption of disjoint task support, which does not hold in the standard image classification merging benchmark. Another hypothesis is related to the linear mode connectivity (Garipov et al., 2018) phenomenon (Frankle et al., 2020; Neyshabur et al., 2020), which concerns the same pretrained model finetuned on the same task with different SGD noise due to hyperparameter difference or data shuffling. The same phenomenon was observed to be layer-wise on modern model architectures (Adilova et al., 2023), and further generalized to cross-task linearity (Zhou et al., 2024b) in the context of model merging, where modes are related to different input tasks. Zhou et al. (2024b) provides a first-order Taylor expansion analysis to prove the existence of cross-task linearity relying on the convexity assumption. Similarly, the convexity assumption also appears in more recent theoretical analyses (Tao et al., 2024) that relates task arithmetic with one-shot federated learning theories, which is not required in our framework. Additionally, Zhou et al. (2024a) assumes mean-squared error as the loss function to analytically compute the optimal vector merging weights. In summary, all existing theories rely on assumptions such as disjoint task support and loss convexity that do not hold in practical models. In contrast, we do not assume loss convexity and further relax the disjoint task support assumption. We provide both discussion and empirical evidence that support our assumptions. For the literature review of other model merging methods, please refer to Appendix B.
Task Grouping
Since it is widely believed that jointly training similar tasks together improves accuracy performance or convergence (Caruana, 1993) in the multi-task learning problem, identifying task groupings is a well-explored area. Classic methods include convex formulations of task clustering (Jacob et al., 2008), latent task basis learning (Kumar & Daume III, 2012), and gradient-based methods (Fifty et al., 2021), which are more suitable for modern deep learning. When task vectors are used for addition in multi-task settings, many techniques from the multitask grouping literature can be adapted to our context. Another interesting recent work proposed the LoRA (Hu et al., 2021) Adapter library (Ostapenko et al., 2024), which uses a similar clustering approach to enable more modular large language models, along with a router-based adapter reuse algorithm. Such advantage was also proved to be successful for LoRA merging (Zhao et al., 2024), which can be seen as a special case under our bases framework. Note that we focus on the broader context of task arithmetic to examine how creating a task vector bases, or library, impacts the performance of multitask, unlearning, and domain generalization capabilities, regardless of fine tuning strategy used for task vectors.
3 Preliminary
Problem Setting
Let $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ be the loss function, and $h:\mathcal{X}\times\Theta\to\mathcal{Y}\subseteq\mathbb{R}$ be the classifier. When the context is clear, we omit some arguments for $\ell$ and $h$. We consider the initial pre-trained model parameter $\theta_{0}\in\mathbb{R}^{d}$, which is fine-tuned on $T$ tasks to yield fine-tuned parameters $\{\theta_{1},\dots,\theta_{T}\}$ with respect to the loss functions $\{\ell_{1},\dots,\ell_{T}\}$. For $n$ training samples drawn from the $i$-th task distribution $\mathcal{D}_{i}$, we denote the population risk evaluated at $\theta$ as $\mathcal{L}_{i}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}[\ell_{i}(h(x,\theta),y)]$.
Task Arithmetic and Applications
Task vectors are defined as $\tau_{i}:=\theta_{i}-\theta_{0},\forall i\in[T]$. Ilharco et al. (2022) discovered two basic arithmetic operations for task vectors. Let $\alpha$ be the scaling coefficient. Task vector addition is achieved by adding weighted task vectors together, $\theta^{*}=\theta_{0}+\sum_{i=1}^{T}\alpha_{i}\tau_{i}$, for multitask learning. Task vector negation involves subtracting a task vector from the pretrained model for machine unlearning, which is $\theta^{*}=\theta_{0}-\alpha\tau$. Built on addition and negation, we can extend to more complex operations for other applications such as domain generalization. In Ilharco et al. (2022), $\alpha$ is tuned on a validation set within 00 to $1$, and all task vectors share the same $\alpha$ to save hyperparameter tuning efforts, although this is not necessarily true for other merging methods.
4 Why Does Task Arithmetic Work?
We will next introduce several practical assumptions based on which we provide our theoretical analysis to explain the benefits of task vectors in model merging. To complement our theoretical analysis, we also provide empirical evidence to verify our theoretical statements in Section 6.
4.1 Assumptions
Assumption 4.1 (Fine-tuning Regime).
We assume that $\forall i\in[T],\frac{\partial\mathcal{L}_{i}(\theta_{i})}{\partial{\theta}}=\mathbf{0}$ and $\exists C>0$ such that $\left\lVert\tau_{i}\right\rVert^{2}\leq C$.
4.1 is often met in practice since $\theta_{i}$ is fine-tuned from the pre-trained model $\theta_{0}$ on the particular downstream task ${\cal D}_{i}$ until convergence. Furthermore, during the fine-tuning regime, the change of model parameters is relatively small. Through a sparsity localization technique, He et al. (2024) show that it is sufficient to only fine-tune 1$\sim$5of the model parameters for competitive performances.
Assumption 4.2 (Task Vector Near Orthogonality).
There exists a universal constant $\epsilon>0$ such that $\forall i\neq j,|\cos(\tau_{i},\tau_{j})|\leq\epsilon$.
A small $\epsilon$ in 4.2 holds when any pair of tasks are not related to each other, which happens when task vectors are restricted to be sparse (He et al., 2024). Cross-task generalization when tasks are similar to each other are better understood (Tripuraneni et al., 2020; Hu et al., 2024).
Remark 4.3.
From a technical perspective, if $\tau_{i}$ and $\tau_{j}$ are independent standard Gaussian random vectors, then $\mathbb{E}[|\cos(\tau_{i},\tau_{j})|]=\sqrt{2/\pi d}$ and $\text{Var}(|\cos(\tau_{i},\tau_{j})|)\approx(1-\frac{2}{\pi})/d$, so $|\cos(\tau_{i},\tau_{j})|\to 0$ when $d\to\infty$ by Chebyshev’s inequality. The upper bound of the task vector norm $C$ is also dependent on the $d$. Since fine-tuning only slightly changes each parameter, let $\tau=(\tau^{1},\cdots,\tau^{d})$, and each entry of the task vector has an $O(1)$ change during the fine-tuning. Then, $\left\lVert\tau\right\rVert^{2}=O(d)$.
Assumption 4.4 (Local Smoothness).
Any fine tuning loss function ${\cal L}$ is $L_{i}$-locally smooth w.r.t. model parameters at $\theta_{i}$, which means for any $\theta\in\Theta$ such that $\|\theta-\theta_{i}\|^{2}=O(C),\mathcal{L}(\theta)-\mathcal{L}(\theta_{i})\leq\left<\theta-\theta_{i},\frac{\partial{\mathcal{L}(\theta_{i})}}{\partial\theta}\right>+\frac{L_{i}}{2}\left\lVert\theta-\theta_{i}\right\rVert^{2}.$ Note that $\theta_{i}$ is the fine-tuned model trained on $\mathcal{D}_{i}$ and $L_{i}=\left\lVert\mathbf{H}(\theta_{i})\right\rVert_{2}$ is the spectral norm of the Hessian matrix of $\mathcal{L}$, evaluated locally at $\theta_{i}$. We hide the subscript of $L_{i}$ when the context is clear.
Assumption 4.5 (Coefficients).
Let $\alpha_{1},\dots,\alpha_{T}$ be the coefficients used to scale the task vector in task arithmetic. We assume $\alpha_{i}\geq 0,\forall i$ and $\sum_{i\in[T]}\alpha_{i}=1$.
4.2 Task Arithmetic Bounds
With assumptions in the previous section, by first-order Taylor expansion w.r.t. the fine tuned model parameter, we can get the following statements for different types of task arithmetic. We defer all proof details to Appendix A.
{restatable}
[Task Addition for Multitask Learning]theoremadd Let task addition $\theta^{*}=\theta_{0}+\sum_{i=1}^{T}\alpha_{i}\tau_{i}$ be the model parameter used for multitask learning, then $\forall i\in[T],\mathcal{L}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})\leq 2L_{i}C(1+\epsilon).$ Section 4.2 shows that as long as the task vectors reside in the fine-tuning regime and task vectors are nearly orthogonal, then a single model obtained by model merging simultaneously performs comparably well on all the tasks. This bound does not depend on the number of tasks $T$. The local smoothness constant $L_{i}$ in the generalization bound implies that a flatter minima is preferred in model merging, which also partially explains the empirical success of the Fisher weighted averaging method (Matena & Raffel, 2022) as $\mathbf{H}$ is also the Fisher information matrix when $\ell$ is the cross-entropy loss which is a log-likelihood.
{restatable}
[Task Negation for Unlearning]theoremnegate Let $\theta_{i}^{*}=\theta_{0}-\alpha_{i}\tau_{i}$ be the model parameter used for unlearning task $i$. Then $\forall j\neq i$, $\mathcal{L}_{j}(\theta_{i}^{*})-\mathcal{L}_{j}(\theta_{0})\leq L_{j}C\left(\frac{3}{2}+\epsilon\right)$. Since $C$ is small due to fine-tuning, $\mathcal{L}_{j}(\theta_{i}^{*})\approx\mathcal{L}_{j}(\theta_{0})$, which means that the negation of a task for forgetting will not adversely impact the performance of other orthogonal tasks, which has been shown empirically in Ilharco et al. (2022), in contrast to other classic unlearning methods like gradient ascent.
{restatable}
[Out-of-Distribution Generalization]theoremgeneralize Given a collection of source task vectors $\mathcal{S}=\{\tau_{1},\tau_{2},\dots,\tau_{T}\}$ and a target task vector with $\|\tau_{\mathrm{tar}}\|^{2}\leq C$. If $\exists i\in[T]$ such that $\langle\tau_{\mathrm{tar}},\tau_{i}\rangle\geq\beta C$ for $0<\beta\leq 1$, then there exists a merging scheme $\alpha_{i},i\in[T]$ such that for the merged model $\theta^{*}=\theta_{0}+\sum_{i=1}^{T}\alpha_{i}\tau_{i},\mathcal{L}_{\mathrm{tar}}(\theta^{*})\leq\mathcal{L}_{\mathrm{tar}}(\theta_{\mathrm{tar}})+L_{\mathrm{tar}}C(1-\beta).$ This implies when $\beta$, which roughly corresponds to the cosine similarity of the two task vectors, is large enough, the gap between $\mathcal{L}_{\mathrm{tar}}(\theta^{*})$ and $\mathcal{L}_{\mathrm{tar}}(\theta_{\mathrm{tar}})$ is small, so we can use the combination of similar task vectors to achieve similar generalization performance for tasks that are out-of-distribution (OOD) w.r.t. the source models.
5 Task Vector Bases
Under limited budget constraints, it is impractical to save all task vectors for a large number of tasks $T$. Although one could argue to only save one final merged model, we lose flexibility of task vector composition, especially when we want to only merge or unlearn a part of the knowledge from a certain task in the future. Besides, the addition performance drop as the number of tasks grows is reported in Ilharco et al. (2022) in practice. Therefore, we propose using task vector bases to reduce the number of task vectors to save while retaining task information maximally and preserving the flexibility of model composition.
Definition 5.1 (Task Vector Bases).
Given $T$ task vectors, task vector bases are a new set of $d$-dimensional vectors $\{\tau_{B_{1}},\dots,\tau_{B_{m}}\}$, where $1<m<T$, created by applying certain transformations to the original $T$ task vectors.
5.1 Bases Creation
From Section 4.2, we want $\epsilon$ to be small, thereby creating near-orthogonal bases. As observed in the green box of Figure 3, non-related tasks are nearly orthogonal, while similar task vectors represent certain interpretable skills. Based on this intuition, we define the transformation in Definition 5.1 by grouping similar task vectors using clustering algorithms to remove redundant information across tasks.
Algorithm 1 Task Vector Bases Creation
1:  Input: task vectors $\tau_{1},\dots,\tau_{T}$, inner merge methods $\texttt{merge}_{\text{in}}()$, threshold $\delta_{1}<0$
2:  $M\in\mathbb{R}^{T\times T},M_{ij}=\cos(\tau_{i},\tau_{j})$ // Task clustering
3:  $k=\arg\max_{1<i\leq{T-1}}\{\lambda^{(i+1)}_{L_{|M|}}-\lambda^{(i)}_{L_{|M|}}\}$
4:  $\mathcal{C}_{1},\dots,\mathcal{C}_{k}=$ spectral_cluster$(|M|,k)$
5:  $\mathcal{C}=\{\}$ // Separate positive and negative directions
6:  for $i\in[k]$ do
7:   Extract submatrix $M_{\mathcal{C}_{i}}$ from $M$
8:    if $\min{M_{\mathcal{C}_{i}}}<\delta_{1}$ then
9:    $M_{\mathcal{C}_{i}}\mathrel{+}=|\min{M_{\mathcal{C}_{i}}}|$
10:    $\mathcal{C}_{pos}^{i},\mathcal{C}_{neg}^{i}=$ spectral_cluster$(M_{\mathcal{C}_{i}},2)$
11:    $\mathcal{C}=\mathcal{C}\cup\{\mathcal{C}_{pos}^{i},\mathcal{C}_{neg}^{i}\}$
12:   else $\mathcal{C}=\mathcal{C}\cup\{\mathcal{C}_{i}\}$
13:  $\mathcal{C}_{B_{1}},\dots,\mathcal{C}_{B_{m}}=\mathcal{C}$
14:  for $i\in[m]$ do
15:   $\tau_{i}=\texttt{merge}_{\text{in}}(\theta_{0},\mathcal{C}_{B_{i}})$ // Create basis by merging
16:  Return: $\{\tau_{1},\dots,\tau_{B_{m}}\}$
Table 1: Bases Arithmetic. Here $i$ is a task id. There are several ways to interpret similarity search given bases and the target task.
Arithmetic Expression
Addition for multitask $\theta^{*}=\texttt{merge}_{\text{out}}(\theta_{0},\{\tau_{1},\dots,\tau_{B_{m}}\})$.
Negation for unlearning $i$ Find $B_{j}$ the most similar to $i$,
$\theta^{*}=\theta_{0}-\alpha\tau_{B_{j}}$.
OOD Generalization on $i$ Find $U=\{B_{j_{1}},\dots,B_{j_{l}}\}$ similar to $i$,
$\theta^{*}=\texttt{merge}_{\text{out}}(\theta_{0},\{\tau_{B_{j_{1}}},\dots,\tau_{B_{j_{l}}}\})$
if $U=\emptyset$, store $\tau_{i}$ as new basis.
In Algorithm 1, we first create the similarity matrix of task vectors and pass this matrix into spectral clustering (Ng et al., 2001) to group the tasks. Next, we can apply any existing merging method as $\texttt{merge}_{\texttt{in}}$ to create task vector bases. When cosine similarity becomes strongly negative, we first use the absolute value of the similarity for clustering and further partition the positive and negative groups with an additional clustering step. However, the scenario where task vectors have nearly opposite directions rarely occurs in practice for natural tasks unless specific adversarial tasks are defined. Therefore, lines 5-9 are typically optional. We choose spectral clustering for memory efficiency. Unlike $k$-means (Lloyd, 1982), which requires storing the entire set of $O(Td)$ task vectors during the update, spectral clustering only requires storing an $O(T^{2})$ matrix for eigenvalue computation, which is more advantageous since $d\gg T$. Additionally, in line 3 of Algorithm 1, we can determine the number of bases by inspecting the eigenspectrum of the Laplacian matrix $L_{|M|}$, without the need for additional hyperparameter tuning. This approach reduces the $T$ task vector addition problem to saving only $m$ task vector bases, significantly reducing the memory footprint.
5.2 Bases Arithmetic
We show how to do arithmetic with task vector bases in Table 1, with only slight modifications from standard task arithmetic operations discussed in Section 4.2. There are two points that need special attention. First, we observe the usage of $\texttt{merge}_{\text{out}}$, which is another iteration of model merging. We can replace $\texttt{merge}_{\text{out}}$ easily with dataless methods such as Model Soup (Wortsman et al., 2022) or TIES (Yadav et al., 2024), while for more advanced methods, we provide an example in Section 6.4.2 to show possible modifications that fit in our framework. Second, there are multiple ways to do the similarity search in Table 1. The naive method is to fine tune $\theta_{i}$ till convergence to get $\tau_{i}$ and simply compute the cosine similarity between $\tau_{i}$ and bases vectors, but then one could argue to directly use $\tau_{i}$ for unlearning and OOD generalization. Fortunately, in Section D.2, we see without training $i$ till convergence, intermediate checkpoints even with one step of gradient update still reflects important task information. Also, as we will see in Section 6.4.1 since bases can be interpreted as some skill: for example, one bases represents digit classification and the target task is MNIST, we can use such heuristics of task similarity knowledge without any explicit computation of $\tau_{i}$. In these two scenarios, since $\tau_{i}$ is either far from the optimal or does not exist, it can be more beneficial to use stored knowledge in the bases.
6 Experiments
6.1 Models and Datasets
This paper includes both computer vision and natural language processing experiments. In the vision experiments, we use CLIP (Radford et al., 2021) and ViT (Dosovitskiy, 2020) models from OpenCLIP (Ilharco et al., 2021), along with the following 8 datasets: MNIST (LeCun et al., 1998), EuroSAT (Helber et al., 2019), Cars (Krause et al., 2013), GTSRB (Stallkamp et al., 2012), RESISC45 (Cheng et al., 2017), DTD (Cimpoi et al., 2014), SUN397 (Xiao et al., 2010), and SVHN (Netzer et al., 2011). The language experiments are based on RoBERTa (Liu, 2019) models, tested on a medium scale benchmark with 12 tasks (Panigrahi et al., 2023; He et al., 2024) and a larger scale collection of 70 tasks, the latter with further details provided in Appendix E.
6.2 Empirical Verification of Theoretical Results
We primarily focus on Section 4.2 and examine the relationship between the loss gap and key constants.
Figure 2: Task vector similarity vs. $\mathcal{L}_{\text{MNIST}}(\theta^{*})-\mathcal{L}_{\text{MNIST}}(\theta_{\text{MNIST}})$, where $\theta^{*}=\theta_{0}+0.5\tau_{\text{MNIST}}+0.5\tau_{\text{task}}$. This figure includes two different set of CLIP ViT/B-32 task vectors. The pink shade includes the high similarity high loss gap region, and the green shade is the low similarity low loss gap region. This implies larger task similarity $\epsilon$ is harmful for addition.
Task Vector Orthogonality $\epsilon$
To verify how task vector similarity $\epsilon$ impacts the performance, we conduct the experiment shown in Figure 2. We merge the MNIST task vector with each of the other task vectors, all having similar norms ranging from $[2,3)$ (see Table 7 in the Appendix for details), and set the scaling coefficient $\alpha$ to $0.5$. In this setting, we approximately control all constants in Section 4.2, including $L$, $\alpha$, and $C$, and observe that highly similar tasks, such as digit classification in MNIST, SVHN, and GTSRB, lead to larger loss gaps or worse performance for MNIST compared to less related tasks.
Figure 3: Task vector similarity matrix for two checkpoints. The left green box represents the task similarity for vectors all derived from fine-tuning pretrained model 1. The right pink box represents the similarity values for task vectors from two different checkpoints, which corresponds to small $\epsilon$ in the 50/50 row of Table 2.
Figure 4: Task vector norms in different settings. Since the distance of two pretrained models ($\approx 314$) are much larger than the distance between the pretrained model and their own fine tuned model ($\approx 2$), in Table 2 if we subtract pretrained 2 from any finetune 1, $\left\lVert\tau\right\rVert$’s upper bound $C$ is huge, leading to the merging failure. For visualization purpose, we only show two randomly selected dimensions, while the numbers for $C_{1},C_{2},\left\lVert\tau\right\rVert$ are directly computed from high-dimensional vectors based on data.
Table 2: Task vector mixing absolute average accuracy, which is the average of all task test performances evaluated with the merged model. Numbers 1 and 2 refer to the identities of the pretrained checkpoints. “50/50” represents the experiment where 50% of the own task vector is mixed with 50% of task vectors derived from the other pretrained model. The task vector $\tau$ is defined as $\theta_{i}-\theta_{0}$.
$\theta_{i}$$\ \theta_{0}$ pretrained 1 pretrained 2
finetune from 1 70.83 51.71
50 / 50 59.92 61.71
finetune from 2 54.21 71.09
Interaction between $C$ and $\epsilon$
We provide additional evidence that both $C$ and $\epsilon$ must be constrained for the success of task vector addition. In Table 2, we collected task vectors for 8 tasks from two CLIP checkpoints pretrained with different hyperparameters. From this table, we observe that successful task addition reveals the identity of the task vectors. For optimal merging performance, we should only add task vectors fine-tuned from the same checkpoint, as any mixture of task vectors from different checkpoints will cause a significant performance drop. The above empirical observation consolidates our 4.1 that task vectors should reside in the same fine-tuning regime. To elaborate, from Figure 3, although all $\epsilon$ values in the pink box are very low, task addition still fails due to the large $C$ value. From Figure 4, we see that with different hyperparameters, the two pretrained models are situated in two local convex basins, and the distance between the two checkpoints is much larger than the task vectors (for more discussion, see Section D.1). Thus, if we create task vectors by subtracting the wrong pretrained checkpoint, the large $C$ value leads to the failure of task addition.
Figure 5: $\mathcal{L}_{\text{DTD}}(\theta^{*})-\mathcal{L}_{\text{DTD}}(\theta_{\text{DTD}})$ by merging $\tau_{\text{DTD}}$ with other task vectors, setting scaling coefficient as $0.5$. Two colored pretrained checkpoints have different local smoothness values.
Local Smoothness $L$
The local smoothness $L$ is specific to each pretrained model due to differences in their optimization trajectories. Since, as shown in Figure 2 and Figure 4, the differences in $C$ and $\epsilon$ (not $\theta_{i}$) between two pretrained models are small. In Figure 5, we merge the DTD task vector with each of the other task vectors and compare the loss gap between two checkpoints. Because it is not feasible to load $\mathbf{H}(\theta_{i})\in\mathbb{R}^{d\times d}$ directly onto the GPU, we estimate $L$ using the power iteration method (Mises & Pollaczek-Geiringer, 1929) to reduce the largest eigenvalue problem to a Hessian-vector product computation. As seen in Figure 5, larger local smoothness consistently leads to a larger gap from the optimal loss term across datasets, resulting in worse merging performance.
6.3 Vision Experiments with Task Vector Bases
Table 3: Addition of MNIST, SVHN, EuroSAT, RESISC45 with Task Arithmetic as $\texttt{merge}_{\text{out}}$ for different $\texttt{merge}_{\text{in}}$. MS-ER represents the partition of (MNIST, SVHN) and (EuroSAT, RESISC45).
$\texttt{merge}_{\text{in}}$ Avg. $k=2$ Avg. All $k$ MS-ER
MTL 95.45 93.85 95.52
TIES 88.68 88.48 90.43
Fisher 82.41 83.67 86.37
Tangent 88.50 88.75 90.05
Topk 82.17 83.05 85.69
Mean 85.65 86.44 88.65
Bases Addition
For a detailed analysis of the behavior of task vector bases methods, we first fix a small subset of tasks where the separation between digit classification (MNIST, SVHN) and satellite image classification (EuroSAT, RESISC45) is obvious and matches the optimal clustering result that reduces the memory storage by half. With only $4$ task vectors, we can enumerate all possible partitions of tasks and understand the position of the merging performance using this natural partition.
In Table 3, we fix the outer merging method $\texttt{merge}_{\text{out}}$ for various inner merging methods $\texttt{merge}_{\text{in}}$, including multi-task joint training (MTL) by mixing all task datasets together during fine-tuning, Fisher merging (Matena & Raffel, 2022), tangent fine-tuning (Ortiz-Jimenez et al., 2024), TIES (Yadav et al., 2024) that resolves the sign conflicts of different task vectors, Topk (similar to the dataless version of He et al. (2024)) that only keeps the top 5% magnitude parameters, and the Mean of all task vectors (Wortsman et al., 2022). We do not report routing-based methods due to the additional introduction of router parameters. These methods represent the most popular categories of model merging. We can observe that MS-ER is consistently better than the average of all $k=2$ and $k\in[1,4]$ partitions, across different merging methods. By grouping, we can achieve better-than-average addition performance while using a small memory budget.
Bases Negation
After the first step of merging in Table 3, due to memory budget, we reduce the number of saved checkpoints. If we want to unlearn part of the knowledge under memory constraints, Table 4 shows the preferred bases for this task. In Table 4, the goal is to unlearn MNIST from the pretrained model, so we tune $\alpha$ to maintain the control metric at 95% of the pretrained model performance, evaluated on ImageNet. Following the setup in Table 3, we save the MS and ER bases with two copies of checkpoints. The alternative extreme is to store only one fully merged model (MSER) from Table 3 to minimize memory usage. However, since MSER mixes unrelated task vectors, it is usually less effective than subtracting a MS basis when the goal is to unlearn MNIST-specific information. For some $\texttt{merge}_{\text{in}}$ methods, such as TIES and Mean, we observe that MS performance is better than directly subtracting the task vector trained on MNIST, likely due to the overlap of skills between MNIST and SVHN. This demonstrates that the memory-efficient task vector bases arithmetic can be highly flexible as well.
Table 4: Unlearning MNIST from pretrained model by negating bases constructed from MNIST, SVHN, EuroSAT, and RESISC45. For Tangent, directly subtracting $\tau_{\text{MNIST}}$ has the target performance of 0.11. For others, this target performance is 15.68.
MS MSER
$\texttt{merge}_{\text{in}}$ Target ($\downarrow$) Control Target ($\downarrow$) Control
MTL 21.53 63.13 31.91 63.60
TIES 12.17 63.18 22.27 62.76
Fisher 17.83 63.49 16.74 63.09
Tangent 0.12 63.00 3.80 63.31
Topk 19.53 64.36 22.89 62.44
Mean 10.36 63.21 20.61 64.27
Table 5: Generalization performance on GTSRB and DTD with different $\texttt{merge}_{\text{in}}$ methods by using bases constructed from MNIST, SVHN, EuroSAT, and RESISC45. The pretrained model performance 36.48 on GTSRB and 54.73 on DTD.
GTSRB DTD
$\texttt{merge}_{\text{in}}$ MS-ER M-S-E-R MS-ER M-S-E-R
MTL 41.02 37.91 54.73 55.00
TIES 43.15 37.91 54.73 55.00
Fisher 41.31 37.91 54.89 55.00
Tangent 39.21 41.59 55.10 55.42
Topk 40.37 37.91 55.00 55.00
Mean 41.97 37.91 54.73 55.00
Bases OOD Generalization
After saving task vector bases from MNIST, SVHN, EuroSAT, and RESISC45 in Section 6.3, we test generalization metrics on unseen tasks in our vision benchmark. We set a similarity threshold to only use the most similar task vector bases for GTSRB and DTD generalization in Table 5, with $\texttt{merge}_{\text{out}}$ set to Task Arithmetic. Since we use only one basis for generalization, this reduces to tuning $\alpha$ for the selected basis on the validation set. For GTSRB, which involves traffic sign recognition, we expect overlap with the digit classification skills in MNIST and SVHN. In Table 5, we observe that the performance of the MS-ER column, corresponding to the bases method, is generally higher than the M-S-E-R column, where full task arithmetic flexibility is retained. For OOD generalization, using all task-specific information related to the target task proves beneficial. However, since DTD is not related to any stored bases, its OOD generalization performance is almost equivalent to the pretrained model. This validates Section 4.2 and shows that when an OOD task is not too far away from seen tasks, leveraging bases knowledge is more effective than relying solely on a task vector trained on a single task, with the added benefit of memory savings.
6.4 Language Experiments with Task Vector Bases
6.4.1 Results for Large Scale Benchmark
Figure 6: Task similarity matrix of merging 70 natural language tasks when $k=3$ (left) and $k=10$ (right), ordered by cluster index based on spectral clustering. We use orange boxes to highlight the clusters with more than one task. When $k=3$, the clusters from top to bottom are: 1-sentence tasks with multiple choice questions (MCQ), 2-sentence tasks, and AI-vs-human text classification. When $k=10$, the white boxes represent natural language inference (NLI) tasks, bAbI (Weston et al., 2015), AI-vs-human, topic classification, two sentiment analysis groups, and MCQ.
Saving Memory Storage with Bases
We collect 70 natural language understanding tasks from Huggingface (Wolf, 2019) datasets that can be reformulated as classification problems. With a larger number of tasks, we can conduct a stress test to determine how much we can compress the memory budget using Algorithm 1. Following Gao et al. (2020); He et al. (2024), we design appropriate prompt templates and labels for masked language modeling fine-tuning of RoBERTa-base, ensuring that each fine-tuned model’s performance exceeds the majority vote baseline. As a result, each of the collected models is meaningful, and since all tasks can be evaluated with accuracy, the merged metrics are directly comparable.
We first examine the result of clustering in Figure 6, where we show the task cosine similarity matrix ordered by the spectral clustering results. We plot two matrices: $k=3$, which represents the number of clusters selected based on the largest eigenvalue gap, and $k=10$, which further partitions tasks into more fine-grained groups. Consequently, The properties of the task vectors change according to the prompt template difference. When $k=3$, the tasks are split based on the number of input sentences, and when $k=10$, there are more fine-grained clusters that reflect the task distribution, such as the separation of bAbI-NLI tasks, which reformulate tasks from Weston et al. (2015) as NLI problems, and other types of NLI tasks. In summary, the generated task clusters in step 2 of Figure 1 are reasonable, as they reflect certain interpretable skills learned by all models. And when we use $k=3$ as the algorithm suggested, we reduce the storage of 70 task vectors into 3 bases, saving 95.71% of the memory budget.
Table 6: Merged accuracy/F1 and storage cost on the $12$-task language benchmark. First two rows contain numbers in He et al. (2024).
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Avg. #Masks SparsityStorage (GB)
Task Arithmetic 88.5 88.2 80.3 82.9 32.0 61.0 62.0 56.1 49.5 65.6 82.8 62.3 67.5 $\times$ 12 100 3.90
Localize-and-Stitch 89.6 89.6 84.9 82.8 78.2 82.0 73.4 62.1 58.0 63.3 82.0 65.1 75.9 $\times$ 12 10.9 0.42
 - Mask Sharing 90.4 89.4 84.8 76.4 89.4 90.9 53.0 58.2 51.9 59.2 77.7 67.3 74.1 $\times$ 4 11.6 0.15
 - Stitch Twice 88.3 88.5 83.5 78.5 78.6 85.1 54.3 74.1 62.8 67.5 81.4 74.2 76.4 $\times$ 12 7.93 0.31
Generalization Performance for Merged Model
Given the task clusters, we now evaluate the performance of using task vector bases. In Figure 7, we plot the performance of the task vector bases against 5 randomly selected partitions of tasks for the same $k$. To avoid the computational overhead of selecting $\alpha$, we set both the inner and outer merging methods to Mean (Wortsman et al., 2022). From the results, we observe that our basis method with $k=3$, selected based on the spectral gap, achieves the best performance overall, outperforming the Mean baseline when $k=1$, as well as the pretrained model and majority vote baselines. Furthermore, we observe a performance degradation of the bases method after $k=3$, and a gradual performance increase for random partitions, although all random partition metrics are worse than the Model Soup baseline. This supports the importance of using the eigengap trick to select $k$ in Algorithm 1, which effectively balances the task addition conflict for two rounds of merging. Based on Figure 7, our bases method surpasses model merging baseline performances in the large $T$ challenging setting.
Figure 7: Absolute average accuracy for each number of bases $k$, comparing the task vector bases method and random partition of tasks. When $k=1$, it reduces to one step of model merging. The pretrained model $\theta_{0}$ performance is $42.15$, and using majority vote for each task’s performance is $40.47$, both of which are far below the two lines in this figure.
6.4.2 Results for Medium Scale Benchmark
Since most merging methods require $\alpha$ tuning or additional training, testing them on the $70$-task benchmark is time-consuming. To better understand how bases methods compare with other popular baselines, we conduct experiments on a medium-scale benchmark with $12$ language tasks.
Bases Addition with Localize-and-Stitch
(L&S) He et al. (2024) divides model merging into two steps: first, localizing the task information by learning a sparse vector mask, and second, stitching sparse models with mask normalization, ensuring that $\alpha$ is a probability vector like 4.5. Compared to other model merging methods, He et al. (2024) offers the dual advantage with significantly improved multitask performance due to reduced conflict from localization, and drastically lowered memory costs by saving only a small proportion of nonzero weights. Thus, to maximally boost all metrics, we investigate how to integrate L&S into our bases creation framework which requires two steps of merging. As described in Section 5.1, after clustering sparse localized task vectors, we can directly apply L&S as $\texttt{merge}_{\text{in}}$ within each cluster. For $\texttt{merge}_{\text{out}}$, it requires special design and we propose two modifications to L&S.
The first approach, Mask Sharing, creates a shared binary mask within each task cluster by averaging the task masks and rounding the averaged mask entries. Formally, if each task $i$ has an associated task-specific mask $\gamma_{i}\in\{0,1\}^{d}$, and the cluster size is $N$, these clustered tasks share one mask $\bar{\gamma}\leftarrow\texttt{round}(\sum_{i}\gamma_{i}/N)$. This approach significantly reduces memory storage, requiring only $m<T$ copies of sparse models, though it may slightly increase mask sparsity. However, this method sacrifices some task-specific information, leading to task conflicts during merging and suboptimal generalization performance in Table 6.
The second approach, Stitch Twice, first localizes tasks and stitches by mask normalization within cluster. Afterward, in $\texttt{merge}_{\text{out}}$, we round within-cluster-normalized masks as updated binary task masks, and use them as input for the second iteration of stitching across all $T$ tasks. Formally, in the first iteration, for $l\in[d]$, the $l$-th entry of the mask for task $i$ is $\gamma_{i}^{l}\leftarrow\texttt{round}(\gamma_{i}^{l}/\sum_{j=1}^{N}\gamma_{j}^{l})$, and the second iteration of mask update is $\gamma_{i}^{l}\leftarrow\gamma_{i}^{l}/\sum_{j=1}^{T}\gamma_{j}^{l}$. Inspired by the improvements shown in Figure 7, this two-step process automatically readjusts task vector weights with appropriate task relationships learned from clustering. Unlike Mask Sharing, Stitch Twice requires $T$ copies of sparse models, since each task retains its own mask. However, in Table 6, Stitch Twice improves generalization performance from L&S (and more competitive baselines, see Section E.1) while also surprisingly reducing memory costs due to increased mask sparsity from double normalization.
In summary, we can integrate L&S into our proposed bases framework with minimal efforts, achieving both enhanced generalization and improved memory efficiency, highlighting the broad applicability of our proposal.
7 Conclusion and Future Work
In this work, we revisited task vectors as a practical approach for model editing. By providing a theoretical foundation for task arithmetic, we addressed key gaps in understanding its underlying mechanisms and identified the conditions for its success. Building on these insights, we introduced a scalable framework that reduces the memory footprint of task vector arithmetic through clustered task bases, significantly improving efficiency while maintaining high performance, and demonstrated the flexibility of bases arithmetic. Looking forward, we foresee several future directions. First, the creation of our task vector bases depends on the task relationship, and there are additional opportunities to improve memory efficiency from hardware perspectives. Additionally, by clustering, we assume that there are disjoint groups of tasks for the initial task vectors. Therefore, it might be interesting to further decompose and reorganize task vectors to boost performance for more complicated task relationships.
Impact Statements
This paper contributes foundational research in the areas of task vector arithmetic and scalable model merging within the machine learning community. Our primary goal is to advance the theoretical understanding and practical methodologies for efficient multi-task learning and model adaptation. Given the scope of this research, we do not anticipate immediate ethical concerns or direct societal consequences. Therefore, we believe there are no specific ethical considerations or immediate societal impacts to be emphasized in the context of this work.
Acknowledgment
SZ, YH, and HZ are partially supported by an NSF IIS grant No. 2416897. HZ would like to thank the support from a Google Research Scholar Award. WY is supported by a gift from AWS AI for research in Trustworthy AI. MY is partly supported by MEXT KAKENHI Grant Number 24K03004 and by JST ASPIRE JPMJAP2302. We are thankful to Wei Hu, Yuki Takezawa, Jose Renato Restom for helpful discussions. The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect the official policies or positions of the supporting companies and government agencies.
Appendix A Proof Details of Section 4.2
[…]
Proof.
Note that since $\theta^{*}=\theta_{0}+\sum_{i=1}^{T}\alpha_{i}\tau_{i}$, so
$\|\theta^{*}-\theta_{0}\|^{2}=\left\|\sum_{i=1}^{T}\alpha_{i}\tau_{i}\right\|^{2}\leq\left(\sum_{i=1}^{T}\alpha_{i}\|\tau_{i}\|\right)^{2}\leq C,$
which means that $\theta^{*}$ is within the fine-tuning regime and satisfies the local smoothness assumption. Hence, if $x\sim\mathcal{D}_{i}$,
${\mathcal{L}}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})$ $\theta^{*}-\theta_{i},\frac{\partial\mathcal{L}_{i}(x,\theta_{i})}{\partial{\theta}}\right>+\frac{L_{i}}{2}\left\lVert\theta^{*}-\theta_{i}\right\rVert^{2}$
[…]
$\leq\sum\limits_{j=1}^{T}\alpha_{j}^{2}\left\lVert\tau_{j}\right\rVert^{2}+2\sum_{j\neq k}\alpha_{j}\alpha_{k}\left|\left<\tau_{j},\tau_{k}\right>\right|+2\sum\limits_{j=1}^{T}\alpha_{j}\left|\left<\tau_{j},\tau_{i}\right>\right|+\left\lVert\tau_{i}\right\rVert^{2}$
$=\sum\limits_{j=1}^{T}\alpha_{j}^{2}\left\lVert\tau_{j}\right\rVert^{2}+2\sum_{j\neq k}\alpha_{j}\alpha_{k}\left|\left<\tau_{j},\tau_{k}\right>\right|+2\sum\limits_{j\neq i}^{T}\alpha_{j}\left|\left<\tau_{j},\tau_{i}\right>\right|+2\alpha_{i}\left\lVert\tau_{i}\right\rVert^{2}+\left\lVert\tau_{i}\right\rVert^{2}$
$\leq\left(\sum\limits_{j=1}^{T}\alpha_{j}^{2}+2\alpha_{i}+1\right)C+2\left(\sum_{j\neq k}\alpha_{j}\alpha_{k}+\sum\limits_{j\neq i}^{T}\alpha_{j}\right)\epsilon C.$
To further upper bound the coefficients of the above two terms, we have
$\sum\limits_{j=1}^{T}\alpha_{j}^{2}+2\alpha_{i}+1\leq\sum_{j=1}^{T}\alpha_{j}+2+1=4,$
and
$42\left(\sum_{j\neq k}\alpha_{j}\alpha_{k}+\sum\limits_{j\neq i}^{T}\alpha_{j}\right)\leq 2\left(\left(\sum_{j=1}^{T}\alpha_{j}\right)\left(\sum_{k=1}^{T}\alpha_{k}\right)+\sum\limits_{j=1}^{T}\alpha_{j}\right)=$ 4 .
To conclude, we have
${\mathcal{L}}_{i}(\theta^{*})-\mathcal{L}_{i}(\theta_{i})\leq\frac{L_{i}C}{2}\left(1+2\alpha_{i}+\sum\limits_{j=1}^{T}\alpha_{j}^{2}\right)+\frac{L_{i}\epsilon C}{2}\left(\sum_{j\neq k}\alpha_{j}\alpha_{k}+\sum\limits_{j\neq i}^{T}\alpha_{j}\right)\leq 2L_{i}C(1+\epsilon).\qed$
[…]
Proof.
First, note that
${\mathcal{L}}_{j}(\theta_{i}^{*})-{\mathcal{L}}_{j}(\theta_{0})\leq{\mathcal{L}}_{j}(\theta_{i}^{*})-{\mathcal{L}}_{j}(\theta_{j})+{\mathcal{L}}_{j}(\theta_{j})-{\mathcal{L}}_{j}(\theta_{0})\leq{\mathcal{L}}_{j}(\theta_{i}^{*})-{\mathcal{L}}_{j}(\theta_{j})+|{\mathcal{L}}_{j}(\theta_{0})-{\mathcal{L}}_{j}(\theta_{j})|$
We will upper bound the last two terms separately. To bound ${\mathcal{L}}_{j}(\theta_{i}^{*})-{\mathcal{L}}_{j}(\theta_{j})$, note that $\|\theta_{i}^{*}-\theta_{j}\|^{2}=\|\alpha_{i}\tau_{i}+\tau_{j}\|^{2}\leq\left(\alpha_{i}\|\tau_{i}\|+\|\tau_{j}\|\right)^{2}\leq 4C$, due to the local smoothness of ${\mathcal{L}}_{j}$ around $\theta_{j}$ and the fact that $\partial{\mathcal{L}}_{j}(\theta_{j})/\partial\theta=\mathbf{0}$, we have
${\mathcal{L}}_{j}(\theta_{i}^{*})-{\mathcal{L}}_{j}(\theta_{j})\leq\frac{L_{j}}{2}\|\theta_{i}^{*}-\theta_{j}\|^{2}=\frac{L_{j}}{2}\|\alpha_{i}\tau_{i}+\tau_{j}\|^{2}\leq\frac{L_{j}}{2}\left(\alpha_{i}^{2}C+2\langle\tau_{i},\tau_{j}\rangle+C\right)\leq L_{j}C(1+\epsilon).$
Similarly, for the second term, we have
$|{\mathcal{L}}_{j}(\theta_{0})-{\mathcal{L}}_{j}(\theta_{j})|\leq\frac{L_{j}}{2}\|\theta_{0}-\theta_{j}\|^{2}\leq\frac{L_{j}C}{2}.$
Combine both above, leading to
${\mathcal{L}}_{j}(\theta_{i}^{*})-{\mathcal{L}}_{j}(\theta_{0})\leq L_{j}C(1+\epsilon)+\frac{L_{j}C}{2}=L_{j}C\left(\frac{3}{2}+\epsilon\right),$
as desired. ∎
[…]
Proof.
Let $i^{*}=\operatorname*{arg\,max}_{i\in[T]}\langle\tau_{\mathrm{tar}},\tau_{i}\rangle$ and choose $\alpha_{i^{*}}=1$, $\alpha_{j}=0,~{}\forall j\neq i^{*}$. Clear $\langle\tau_{\mathrm{tar}},\tau_{i^{*}}\rangle\geq\beta C$ and $\theta^{*}=\theta_{0}+\tau_{i^{*}}$. It is easy to check that $\|\theta^{*}-\theta_{\mathrm{tar}}\|\leq 4C$ so by the local smoothness assumption of ${\mathcal{L}}_{\mathrm{tar}}$, we have
${\mathcal{L}}_{\mathrm{tar}}(\theta^{*})-{\mathcal{L}}_{\mathrm{tar}}(\theta_{\mathrm{tar}})$ $\leq\frac{L_{\mathrm{tar}}}{2}\|\theta^{*}-\theta_{\mathrm{tar}}\|^{2}$
$=\frac{L_{\mathrm{tar}}}{2}\|\tau_{i^{*}}-\tau_{\mathrm{tar}}\|^{2}$
$\leq\frac{L_{\mathrm{tar}}}{2}\left(C-2\langle\tau_{\mathrm{tar}},\tau_{i^{*}}\rangle+C\right)$
$\leq L_{\mathrm{tar}}C(1-\beta).\qed$
Appendix B Additional Related Work
B.1 Single-task Merging Methods
Prior to Task Arithmetic (Ilharco et al., 2022), researchers discussed how to combine models fine-tuned on the same task, with some minor differences due to hyperparameter changes, as an alternative to ensembles, starting with model soup (Wortsman et al., 2022). Since fine-tuned models capture more domain-specific skills while pretrained models contain more generic knowledge, WiSE-FT (Wortsman et al., 2021) proposed merging the pretrained model and the fine-tuned model via linear interpolation, achieving balanced or even optimal performance on both in-domain and out-of-distribution generalization metrics. Izmailov et al. (2018) introduced stochastic weight averaging, which includes intermediate checkpoints before model convergence for model merging. Several close variants, such as exponentially moving averaging (Szegedy et al., 2016) and LAtest Weight Averaging (Kaddour, 2022; Sanyal et al., 2023), have been explained theoretically under a unified framework (Wang et al., 2024b).
B.2 Multi-task Merging Methods
The major difference from Section B.1 is that all methods discussed in this subsection focus on the setting that one pretrained model is fine tuned on many different tasks. Task arithmetic (Ilharco et al., 2022) can be seen as the generalization of the single-task model merging method, model soup (Wortsman et al., 2022), where task vectors are simply averaged. In Ilharco et al. (2022), however, the scaling coefficients $\alpha$ are allowed to be tuned. Since then, several ideas have been proposed to improve task arithmetic. First, since tuning $\alpha$ is time-consuming, popular approaches such as Fisher merging (Matena & Raffel, 2022), RegMean (Jin et al., 2022), AdaMerging (Yang et al., 2023), Evol (Akiba et al., 2024) aim to find better methods to automatically adjust scaling coefficients for improved task arithmetic performance. Second, instead of using standard fine-tuning to obtain $\tau$, alternative fine-tuning methods, such as tangent space fine-tuning (Ortiz-Jimenez et al., 2024) and parameter-efficient fine-tuning methods (Zhang et al., 2023; Tang et al., 2023b; Stoica et al., 2024), are employed in task arithmetic to disentangle task information for better merging. Third, to reduce task vector conflicts, task vectors can be sparsified into different subspaces by localization (He et al., 2024; Yadav et al., 2024; Yu et al., 2024; Davari & Belilovsky, 2025; Tang et al., 2023a; Wang et al., 2024a), i.e., by masking out useless parameters for each task. Note this category of work has the additional benefit on memory efficiency due to the storage of sparse model weights. Finally, inspired by the Mixture-of-Experts (Shazeer et al., 2017) mechanism, task vector merging performance can be enhanced by learned routers that dynamically merge task-specific and task-shared information (Lu et al., 2024; Tang et al., 2024). For more details on the latest task arithmetic methods and their applications, we refer readers to the model merging survey (Yang et al., 2024).
Appendix C Illustration of Memory Efficiency of Task Vector Bases
In Figure 8, we illustrate the memory advantage of using task vector bases methods, from the perspective of the OOD generalization application. First, in the left figure, we observe that all storage space is used up with $T$ task vectors before clustering. In the second step, since using $m$ bases provides comparable accuracy performance to using $T$ vectors, we only need to store $m$ copies of the checkpoint. When new tasks are very different from the existing memory, as shown in Table 1, we allow $T-m$ new bases to store additional information. Thus, by reducing redundant task vectors to task vector bases, we save space for storing new vectors during domain generalization. This approach is especially beneficial when $T$ original task vectors are highly similar, and future target tasks are orthogonal to the old knowledge.
Figure 8: Workflow of using task vector bases for generation. The light gray box represents all available memory and we assume $T$ task vectors uses up the full memory as shown in the left. Colorful arrows represents different task vector bases. We use clustering and merging from left to middle figure for the bases creation step in Algorithm 1, and use the OOD generalization rule in Table 1 from middle to right which involves saving new bases into the memory. If we simply use the left figure for OOD generalization, we do not allow additional new bases which can lead to trivial performance based on Table 12.
Appendix D Details for Computer Vision Experiments
D.1 Task Vector Norm $C$
Model Dataset RatioNorm Task Vector
laion2b_e16 MNIST 0.46 2.18
EuroSAT 0.45 2.16
Cars 0.54 2.52
DTD 0.39 1.81
GTSRB 0.49 2.32
RESISC45 0.54 2.55
SUN397 0.65 3.03
SVHN 0.56 2.64
laion2b_s34b_b79k MNIST 0.42 2.30
EuroSAT 0.42 2.27
Cars 0.48 2.59
DTD 0.33 1.79
GTSRB 0.45 2.44
RESISC45 0.48 2.63
SUN397 0.58 3.18
SVHN 0.51 2.76
Table 7: Ratio and Norm Task Vector for Different Models and Datasets
Two openclip checkpoints are details can be found from https://github.com/mlfoundations/open_clip/blob/main/docs/PRETRAINED.md where laion2b_s34b_b79k is reported to be trained with larger batch size and learning rate, while two models share the same training data LAION-2B (Schuhmann et al., 2022). In Table 7, we reported the task vector norm and the ratio of task vector norm over the pretrained model norm, which is very small across datasets and models.
We elaborate on the connection of small task vector norm $C$ requirement with previous literature. Ilharco et al. (2022) in its Figure 7 demonstrates that the performance of merging task vectors derived from intermediate checkpoints, far before model convergence, is close to the performance of merging converged task vectors. These intermediate checkpoints typically have smaller norms due to fewer optimization steps, so a small $C$ appears sufficient for the success of task addition. On the other hand, Figure 6 of Ilharco et al. (2022) also indicated that a smaller learning rate is more important for task addition than for standard single-task fine-tuning, which implies that a smaller $C$ is also necessary.
D.2 Identifying Task Groups with Less Training
Step 0 Refer to caption
Step 1 Refer to caption
Epoch 0 Refer to caption
Epoch 1 Refer to caption
Figure 9: Evolution of the similarity matrix for eight vision tasks during fine-tuning, showing the first two steps of batch updates and two full epochs for each task.
When we do not have access to all single-task task vectors at the beginning and the number of tasks $T$ is large, the most time-consuming step in generating the task similarity matrix, as shown in Figure 6, is fine-tuning on each task. In Figure 9, we observe how the task similarity matrix evolves as we increase the number of training steps on an 8-task standard vision benchmark. If we apply spectral clustering to all four matrices—Step 0, Step 1, and Epoch 0—share the same clustering result: [(MNIST, SVHN), (SUN397, DTD), (EuroSAT, RESISC45), (Cars), (GTSRB)]. Starting from Epoch 1, the clusters converge to the same result: [(MNIST, SVHN, GTSRB), (DTD), (EuroSAT, RESISC45), (SUN397), (Cars)]. All matrices in Figure 9 exhibit similar patterns, especially for the two major task groups: digit classification for MNIST, SVHN, and satellite image classification for EuroSAT and RESISC45. Over time, the model learns new skills, such as how traffic sign recognition in GTSRB involves digit identification, similar to MNIST and SVHN. However, these intermediate checkpoints, even with just one batch or one epoch update, are able to reflect useful information about task vectors for creating task bases, eliminating the need for converged task models in Table 1.
D.3 Bases Addition
We report the full table of Table 3 in Table 8 for all possible partitions of MNIST, SVHN, EuroSAT, and RESISC45. We can draw several conclusions. First, our natural partition is almost always among the best for a 2-2 partition. In the context of multitask learning, the common wisdom is that similar tasks can be trained to leverage mutually shared information for better performance (Caruana, 1993). Additionally, by Section 4.2, at the second $\texttt{merge}_{\text{out}}$ step, it is also beneficial to combine two task vector bases with fewer conflicts or smaller $\epsilon$. These two conditions are automatically satisfied by the clustering process. Second, MS-ER is consistently better than the average of all $k=2$ and $k\in[1,4]$ partitions. By grouping, we can achieve better-than-average addition performance while using a small memory budget. Note that there is no consistent partition that always achieves the best performance for all $\texttt{merge}_{\text{in}}$ methods.
In Table 9 and Table 10, we cannot directly use non-dataless merging methods for the outer merging method $\texttt{merge}_{\text{out}}$ as we pointed out in Section 5.2. For example, after merging MS at step one, it is unnecessary to jointly fine-tune MS from scratch at step two or recompute the same Fisher-based weights again. For the remaining dataless methods, we report the merging performance for different $\texttt{merge}_{\text{out}}$ methods, while fixing Mean as $\texttt{merge}_{\text{in}}$ in Table 9, and fixing both steps of merging methods across different sizes of CLIP architectures in Table 10. We observe some similar patterns as in Table 8 across merging methods and architectures. Note that when the task number is $4$, if we use the merging algorithm Mean twice, task vector weight when $k=1$ will be the same as all other 2-2 partitions. Therefore, we see the same numbers for Mean rows in these situations. Although in both tables, MS-ER is not always the best clustering among all 2-2 partitions, we can still see that the MS-ER natural partition is better than average metrics. Finally, by comparing each row, we can say the absolute task vector bases method performance depends on the effectiveness of chosen merging methods. This is more obvious in Table 8 where joint fine tuning methods including MTL and Tangent surpass other basic methods like Mean and Topk by a large margin.
Table 8: Merging MNIST, SVHN, EuroSAT, RESISC45 fixing Task Arithmetic as outer merging method $\texttt{merge}_{\text{out}}$ for different inner merging methods $\texttt{merge}_{\text{in}}$. We included all possible partitions in this table, where each dataset is represented by its first letter, and the partition can be reflected by the location of -. For example, cluster algorithms produces the grouping of MS-ER. When $k=1$, it reduces to one step of merging methods listed at the leftmost column of $\texttt{merge}_{\text{in}}$, and when $k=4$, it reduces to one step of $\texttt{merge}_{\text{out}}$ applied to all task vectors. We bold the natural clustering MS-ER if this cell value of Abs. Avg. Acc. is greater than both Avg. $k=2$ and $k\in[1,4]$.
Avg. $k=1$ $k=2$ $k=3$ $k=4$
$\texttt{merge}_{\text{in}}$ $k=2$ $k\in[1,4]$ MSER MS-ER MR-ES ME-SR M-ESR S-MER E-MSR R-MSE M-S-ER S-R-EM M-R-ES E-S-RM M-E-RS E-R-MS M-S-E-R
MTL 95.45 93.85 98.05 95.52 95.27 94.40 97.26 95.99 96.60 93.12 91.74 90.36 92.25 92.81 93.38 92.42 88.62
TIES 88.68 88.48 90.74 90.43 90.17 90.36 83.69 86.26 89.70 90.17 83.27 88.83 86.32 90.10 88.00 90.58 88.62
Fisher 82.41 83.67 72.88 86.37 81.45 84.21 73.60 83.92 85.72 81.62 82.33 88.36 79.63 90.51 85.78 90.09 88.62
Tangent 88.50 88.75 88.61 90.05 90.05 90.05 85.07 90.36 83.56 90.38 89.39 92.15 88.75 88.57 85.23 89.00 90.01
Topk 82.17 83.05 85.87 85.69 86.16 85.95 75.50 78.55 81.20 82.14 76.43 84.49 79.97 87.83 84.01 83.34 88.62
Mean 85.65 86.44 87.41 88.65 88.65 88.65 79.54 83.13 84.26 86.68 83.14 88.34 85.39 88.75 86.25 89.14 88.62
Table 9: Merging MNIST, SVHN, EuroSAT, RESISC45 fixing model soup as inner merging method $\texttt{merge}_{\text{in}}$ for different outer merging methods $\texttt{merge}_{\text{out}}$.
Avg. $k=1$ $k=2$ $k=3$ $k=4$
$\texttt{merge}_{\text{out}}$ $k=2$ $k\in[1,4]$ MSER MS-ER MR-ES ME-SR M-ESR S-MER E-MSR R-MSE M-S-ER S-R-EM M-R-ES E-S-RM M-E-RS E-R-MS M-S-E-R
Mean 84.74 85.56 87.42 87.42 87.42 87.42 79.50 83.13 84.10 84.22 83.12 86.61 84.22 88.18 85.91 87.25 87.42
TIES 85.57 86.48 83.46 89.60 89.26 89.02 79.86 82.50 84.56 84.18 83.04 87.67 85.61 89.48 87.34 90.90 90.70
Topk 68.24 68.13 68.71 69.01 68.12 67.96 68.71 69.62 67.63 66.65 69.14 67.88 67.60 68.23 67.82 67.09 67.84
Table 10: Merging MNIST, SVHN, EuroSAT, RESISC45 with the same $\texttt{merge}_{\text{in}}$ and $\texttt{merge}_{\text{out}}$ across various CLIP architectures.
Avg. $k=1$ $k=2$ $k=3$ $k=4$
ViT/B-16 $k=2$ $k\in[1,4]$ MSER MS-ER MR-ES ME-SR M-ESR S-MER E-MSR R-MSE M-S-ER S-R-EM M-R-ES E-S-RM M-E-RS E-R-MS M-S-E-R
Mean 86.76 87.48 89.23 89.23 89.23 89.23 82.97 85.41 84.25 86.99 85.57 89.25 87.05 89.01 87.10 88.39 89.23
TIES 84.14 86.21 89.23 90.65 90.73 90.83 78.09 75.97 79.88 82.82 81.35 89.28 87.66 88.85 86.55 89.75 91.55
Topk 71.82 71.57 72.98 72.50 72.26 72.14 71.98 71.62 72.03 70.21 71.45 70.47 70.81 71.56 71.86 70.96 70.75
ViT/B-32 $k=2$ $k\in[1,4]$ MSER MS-ER MR-ES ME-SR M-ESR S-MER E-MSR R-MSE M-S-ER S-R-EM M-R-ES E-S-RM M-E-RS E-R-MS M-S-E-R
Mean 84.74 85.56 87.42 87.42 87.42 87.42 79.50 83.13 84.10 84.22 83.12 86.61 84.22 88.18 85.91 87.25 87.42
TIES 88.24 88.81 85.31 89.28 89.35 90.32 82.90 86.26 91.36 88.19 85.75 89.92 87.64 91.96 90.47 92.78 90.70
Topk 68.86 68.57 70.60 68.56 69.23 69.37 69.62 69.69 68.33 67.21 67.84 67.96 68.04 68.36 68.59 67.04 68.04
ViT/L-14 $k=2$ $k\in[1,4]$ MSER MS-ER MR-ES ME-SR M-ESR S-MER E-MSR R-MSE M-S-ER S-R-EM M-R-ES E-S-RM M-E-RS E-R-MS M-S-E-R
Mean 91.56 91.92 92.98 92.98 92.98 92.98 88.75 92.40 89.11 91.70 91.57 93.94 91.41 92.84 90.19 92.06 92.98
TIES 90.35 91.63 92.98 94.17 94.15 94.67 86.82 88.61 85.88 88.12 91.55 94.53 91.51 93.26 90.28 92.81 95.18
Topk 79.36 79.33 80.02 79.88 79.40 79.48 78.87 79.21 80.12 78.56 78.97 78.57 78.81 79.45 79.46 80.12 79.05
D.4 Bases OOD Generalization
Figure 10: Average cosine similarity of held-out task vectors with the most similar task vector basis for all possible MSER partitions in the setting of Table 3.
We report the cosine similarity of the most similar task vector bases among all 15 MSER partitions for each dataset and each $\texttt{merge}_{\text{in}}$ method in Figure 10. Following our intuition, for most merging methods, GTSRB has the highest similarity due to the overlap between traffic sign identification and digit classification. In contrast, other datasets such as Cars, DTD, and SUN397 do not have a strong relationship with the stored task vector bases. Based on Table 1, it is more recommended to create new task vector bases for each of these datasets.
We include the expanded version of Table 5 in Table 11 and Table 12. When the target task is GTSRB in Table 11, we can observe several patterns. First, in most cases, GTSRB identifies bases containing at least one of MNIST or SVHN. Second, we observe that applying another merging method $\texttt{merge}_{\text{out}}$ is necessary. In our case, we use task arithmetic and report the performance with and without tuning $\alpha$. Without weight tuning, the generalization performance can be worse than the pretrained model performance. Finally, we also added the MSER column, where only 1 final merged checkpoint is stored in memory. Compared to using bases, we can further improve domain generalization performance by using this well-combined basis that contains all old task information. However, due to the significant loss of unlearning capability shown in Table 4, we do not recommend to only save MSER as basis. The equivalent solution for improved OOD generalization performance is to further merge MS and ER only in the OOD generalization step.
In Table 12, however, when we replace the target task with DTD, we do not observe significant improvement using any of the MSER partitions. This is expected because, first, based on the clustering results in Section D.2, DTD is not grouped with any of the MSER tasks, and second, in Figure 10, DTD’s cosine similarity score is consistently low.
Table 11: Detailed version of Table 5 when the target task is GTSRB.
$\texttt{merge}_{\text{in}}$
Success Rate
of finding M/S
Avg.
$\alpha=1$
Avg.
$\alpha$ tuned
MS-ER
$\alpha=1$
MS-ER
$\alpha$ tuned
M-S-E-R
$\alpha=1$
M-S-E-R
$\alpha$ tuned
MSER
$\alpha=1$
MSER
$\alpha$ tuned
MTL 14/15 28.72 43.21 23.68 41.02 22.69 37.91 30.61 44.77
TIES 15/15 33.43 41.16 38.85 43.15 22.69 37.91 42.94 45.59
Fisher 15/15 34.32 40.76 36.08 41.31 22.69 37.91 38.35 41.80
Tangent 14/15 40.35 40.65 38.93 39.21 40.67 41.59 41.77 41.77
Topk 15/15 33.82 40.14 40.37 40.37 22.69 37.91 43.68 43.68
Mean 15/15 38.79 41.85 38.17 41.97 22.69 37.91 44.00 44.13
Table 12: Detailed version of Table 5 when the target task is DTD.
$\texttt{merge}_{\text{in}}$
Avg.
$\alpha=1$
Avg.
$\alpha$ tuned
MS-ER
$\alpha=1$
MS-ER
$\alpha$ tuned
M-S-E-R
$\alpha=1$
M-S-E-R
$\alpha$ tuned
MSER
$\alpha=1$
MSER
$\alpha$ tuned
MTL 33.13 54.75 38.61 54.73 46.27 55.00 27.18 54.73
TIES 46.08 54.82 48.35 54.73 46.27 55.00 45.85 54.73
Fisher 46.86 54.84 48.08 54.89 46.27 55.00 42.65 54.73
Tangent 52.94 55.08 51.70 55.10 50.85 55.42 41.77 41.77
Topk 48.72 54.89 54.09 55.00 46.27 55.00 53.61 54.78
Mean 49.09 54.75 48.19 54.73 46.27 55.00 50.53 54.73
D.5 Model Merging Methods Settings
Figure 11: Cosine similarity matrix when we use tangent linearized fine tuning.
For all model merging methods in Table 3, we follow the exact same hyperparameters as in the corresponding model merging method literature. We add some additional comments below.
Multitask Learning
We keep all the fine-tuning hyperparameters the same as in Task Arithmetic (Ilharco et al., 2022), with the only difference being that when we jointly train on more than one dataset, the total number of training epochs is the sum of all epochs used in Ilharco et al. (2022) for training independently on each dataset. Therefore, the total training time spent on all data points in the 8 datasets is the same for both independent and multi-task training.
TIES
We uses the dataless version of TIES (Yadav et al., 2024), so $\alpha$ is always set to be $0.4$ as recommended.
Tangent
For all tangent method rows in our tables, we jointly tune the corresponding datasets in one cluster, similar to MTL, and replace the standard non-linear fine-tuning with the linearized fine-tuning methods in (Ortiz-Jimenez et al., 2024). For fine-tuning-based methods in model merging, we note that they can still be used for clustering. For example, when we use Figure 11 as the affinity matrix input, the clustering result is: [(MNIST, SVHN), (SUN397, DTD, Cars), (EuroSAT), (RESISC45), (GTSRB)], where we can still observe alignment with both intermediate clustering results and the final converged clustering results in Section D.2. Additionally, the same similarity metric is used for LoRA modules in other contexts, as seen in Ostapenko et al. (2024).
Topk
is inspired from the dataless version of He et al. (2024) by only keeping the top 5% magnitude of the vector entries. To merge the sparse vectors, the update rule of the mask for task $i$ is $\gamma_{i}\leftarrow\gamma_{i}/T$.
Appendix E Details for Language Experiments
E.1 Merging Performance on the Medium-Scale Benchmark
We include the full version of Table 6 in Section E.1. Since we employ a strong merge algorithm, Localize-and-Stitch (L&S), within our framework, we observe the advantage of the two bases L&S versions over several important baselines, including Task Arithmetic, TIES (Yadav et al., 2024), Fisher merging (Matena & Raffel, 2022), RegMean (Jin et al., 2022), AdaMerging (Yang et al., 2023), and Consensus methods (Wang et al., 2024a). Even though Mask Sharing achieves the lowest performance among all L&S variants due to its focus on memory efficiency, its multitask performance still outperforms all other non-L&S baselines.
Table 13: Model merging generalization comparison full table expanded on Table 6, except for last two shaded rows, all numbers are from He et al. (2024). We only report data-based merging algorithms for the best merging performance.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Avg.
Task Arithmetic 88.5 88.2 80.3 82.9 32.0 61.0 62.0 56.1 49.5 65.6 82.8 62.3 67.5
TIES 88.6 88.0 85.2 83.5 22.6 48.2 54.8 35.9 39.7 59.4 79.4 60.3 62.1
Fisher merging 90.0 89.8 83.7 75.8 26.0 54.6 54.2 72.5 65.2 65.6 83.3 67.7 69.0
RegMean 89.7 89.7 84.7 82.6 73.0 79.1 55.9 68.3 56.8 63.8 79.4 64.2 73.9
AdaMerging 85.0 86.1 77.8 81.5 23.0 59.5 61.2 54.1 40.4 54.7 82.2 58.8 63.7
Consensus TA 89.2 89.4 86.6 88.2 37.6 59.6 71.4 66.5 52.0 63.0 88.1 66.4 71.5
Consensus TIES 89.8 89.3 84.7 86.2 31.4 63.1 68.9 60.4 46.1 63.0 86.2 64.4 69.5
Localize-and-Stitch 89.6 89.6 84.9 82.8 78.2 82.0 73.4 62.1 58.0 63.3 82.0 65.1 75.9
 - Mask Sharing 90.4 89.4 84.8 76.4 89.4 90.9 53.0 58.2 51.9 59.2 77.7 67.3 74.1
 - Stitch Twice 88.3 88.5 83.5 78.5 78.6 85.1 54.3 74.1 62.8 67.5 81.4 74.2 76.4
E.2 Hyperparameters
We follow the fine-tuning setting of Gao et al. (2020) by reformulating a 64-shot classification problem into a masked language modeling problem with appropriate templates and prompts.
During fine-tuning, we fix the language modeling head and all embedding layers. For other parameters, we use the AdamW (Loshchilov, 2017) optimizer and a linear learning rate scheduler with a learning rate of $2\times 10^{-5}$, training for 10 epochs with a batch size of 4. For each class label in the dataset, we randomly sample 64 data points from the training set and evaluate the model on the original test set if the label exists. If not, we use the validation set for evaluation. The maximum sequence length is 512 tokens, and unknown text labels are added as new tokens. We run all experiments on NVIDIA RTX A6000 GPUs with 48GB memory.
E.3 Language Tasks in the Large-Scale Benchmark
Now we list all datasets selected after filtering out those with performance lower than the majority baseline.
QNLI
(Wang, 2018) is a question-answering inference task. The prompt template is: <S1>? [MASK] , <S2>. The label mapping is: entailment: Yes, not_entailment: No. The Huggingface link is: https://huggingface.co/datasets/nyu-mll/glue.
MRPC
Microsoft Research Paraphrase Corpus (Dolan & Brockett, 2005) is a task to determine whether the input pair of sentences are semantically equivalent. The prompt template is: <S1> [MASK] , <S2>. The label mapping is: similar: Yes, not_similar: No. The Huggingface link is: https://huggingface.co/datasets/nyu-mll/glue.
CoLA
The Corpus of Linguistic Acceptability (Warstadt, 2019) is the task to determine whether the input sentence is grammatically correct. The prompt template is: <S1> This is [MASK]. The label mapping is: acceptable: Yes, not_acceptable: No. The Huggingface link is: https://huggingface.co/datasets/nyu-mll/glue.
MNLI
MultiNLI (Williams et al., 2017) is a natural language inference task. We use the matched subset for evaluation. The prompt template is: <S1>? [MASK] , <S2>. The label mapping is: entailment: Yes, neutral: Maybe, not_entailment: No. The Huggingface link is: https://huggingface.co/datasets/nyu-mll/glue.
RTE
The Recognizing Textual Entailment (Wang, 2018) dataset is a natural language inference task. The prompt template is: <S1>? [MASK] , <S2>. The label mapping is: entailment: Yes, not_entailment: No. The Huggingface link is: https://huggingface.co/datasets/nyu-mll/glue.
SST2
Stanford Sentiment Treebank-2 (Socher et al., 2013) is a sentiment analysis task with binary labels. The prompt template is: <S1> It was [MASK]. The label mapping is: positive: great, negative: terrible. The Huggingface link is: https://huggingface.co/datasets/nyu-mll/glue.
SST5
is the 5-label version of Stanford Sentiment Treebank (Socher et al., 2013) sentiment analysis dataset. The prompt template is: <S1> It was [MASK]. The label mapping is: positive: great, somewhat_positive: good, neutral: okay, somewhat_negative: bad, negative: terrible. The Huggingface link is: https://huggingface.co/datasets/SetFit/sst5.
SNLI
Stanford Natural Language Inference (Bowman et al., 2015) is a natural language inference task. The prompt template is: <S1>? [MASK] , <S2>. The label mapping is: entailment: Yes, neutral: Maybe, not_entailment: No. The Huggingface link is: https://huggingface.co/datasets/stanfordnlp/snli.
TREC
Text REtrieval Conference (Li & Roth, 2002) is a text classification task that classifies question into $6$ categories. The prompt template is: [MASK]: <S1>. The label mapping is: ABBR: Expression, ENTY: Entity, DESC: Description, HUM: Human, LOC: Location, NUM: number. The Huggingface link is: https://huggingface.co/datasets/CogComp/trec.
SUBJ
Subjectivity (Pang & Lee, 2004) dataset is a task to determine whether given text is subjective. The prompt template is: <S1> This is [MASK]. The label mapping is: subjective: subjective, objective: objective. The dataset link is: http://www.cs.cornell.edu/people/pabo/movie-review-data/rotten_imdb.tar.gz.
CR
Customer Review (Hu & Liu, 2004) is a binary sentiment analysis task. The prompt template is: <S1> It was [MASK]. The label mapping is: positive: great, negative: terrible. The dataset link is https://nlp.cs.princeton.edu/projects/lm-bff/datasets.tar.
MPQA
Multi-Perspective Question Answering (Wiebe et al., 2005) is a binary sentiment analysis task. The prompt template is: <S1> It was [MASK]. The label mapping is: positive: great, negative: terrible. The dataset link is https://nlp.cs.princeton.edu/projects/lm-bff/datasets.tar.
MR
Movie Reviews (Pang & Lee, 2004) is a binary sentiment analysis task. The prompt template is: <S1> It was [MASK]. The label mapping is: positive: great, negative: terrible. The dataset link is https://nlp.cs.princeton.edu/projects/lm-bff/datasets.tar.
AG News
(Zhang et al., 2015) is a topic classification task that classifies news article into 4 categories. The prompt template is: <S1> This is about [MASK] news. The label mapping is: World: international, Sports: sports, Business: business, Sci/Tech: science. The Huggingface link is: https://huggingface.co/datasets/fancyzhx/ag_news.
Yelp
(Zhang et al., 2015) contains yelp reviews and is a sentiment analysis task. The prompt template is: <S1> This place is [MASK]. The label mapping is: 1-star: poor, 2-star: fair, 3-star: good, 4-star: great, 5-star: excellent. The Huggingface link is: https://huggingface.co/datasets/Yelp/yelp_review_full.
IMDb
(Maas et al., 2011) is a movie review binary sentiment analysis dataset. The prompt template is: <S1> This movie is [MASK]. The label mapping is: positive: great, negative: terrible. The Huggingface link is https://huggingface.co/datasets/stanfordnlp/imdb.
Yahoo! Answers
(Zhang et al., 2015) is a question-answer topic classification dataset. The prompt template is: <S1> This is related to [MASK]. The question title, question content, and the best answer are concatenated to be S1. The label mapping is: Society_&_Culture : society, Science_&_Mathematics: science, Health: health, Education_&_Reference: education, Computers_&_Internet: computer, Sports: sports, Business_&_Finance: finance, Entertainment_&_Music: entertainment, Family_&_Relationships: relationship, Politics_&_Government: government. The Huggingface link is https://huggingface.co/datasets/community-datasets/yahoo_answers_topics.
ANLI
(Nie et al., 2019) are three natural language inference tasks collected iteratively and adversarially. The prompt template is: <S1>? [MASK] , <S2>. The label mapping is: entailment: Yes, neutral: Maybe, not_entailment: No. The Huggingface link is: https://huggingface.co/datasets/facebook/anli.
CB
CommitmentBank (De Marneffe et al., 2019) is a natural language inference task. The prompt template is: <S1>? [MASK] , <S2>. The label mapping is: entailment: Yes, neutral: Maybe, not_entailment: No. The Huggingface link is: https://huggingface.co/datasets/aps/super_glue.
WiC
Word-in-Context (Pilehvar & Camacho-Collados, 2018) is a binary classification task to answer whether the same word is used in the same way in two sentences. The prompt template is: <S1> <S2> Does <WORD> have the same meaning in both sentences? [MASK]. The label mapping is: Yes: Yes, No: No. The Huggingface link is: https://huggingface.co/datasets/aps/super_glue.
ETHICS Commonsense
is the commonsense subset of the ETHICS (Hendrycks et al., 2020) dataset, which is a binary classification task to determine whether the behavior of the text matches commonsense moral. The prompt template is: <S1> It was [MASK]. The label mapping is: 0: acceptable, 1: unacceptable. The Huggingface link is: https://huggingface.co/datasets/hendrycks/ethics.
ETHICS Deontology
is the deontology subset of the ETHICS (Hendrycks et al., 2020) dataset, which is a binary classification task to determine whether the excuse is reasonable given the scenario. The prompt template is: <S1> <S2> This is a [MASK] excuse. The label mapping is: 0: great, 1: terrible. The Huggingface link is: https://huggingface.co/datasets/hendrycks/ethics.
ETHICS Justice
is the justice subset of the ETHICS (Hendrycks et al., 2020) dataset, which is a binary classification task to determine whether the text follows the principle of justice. The prompt template is: <S1> It was [MASK]. The label mapping is: 0: unfair, 1: fair. The Huggingface link is: https://huggingface.co/datasets/hendrycks/ethics.
Logiqa2.0 NLI
(Liu et al., 2023) is a set of natural language inference tasks specifically designed for evaluating logical reasoning. The prompt template is: <S1>? [MASK] , <S2>. S1 is the concatenation of major and minor premise. The label mapping is: entailment: Yes, not_entailment: No. The Huggingface link is: https://huggingface.co/datasets/baber/logiqa2.
Amazon Reviews
is a collection of Amazon reviews with rating labels. The prompt template is: <S1> This product is [MASK]. The label mapping is: 1-star: poor, 2-star: fair, 3-star: good, 4-star: great, 5-star: excellent. The Huggingface link is: https://huggingface.co/datasets/SetFit/amazon_reviews_multi_en.
TweetEval Emotion
(Mohammad et al., 2018) is a subset of TweetEval (Barbieri et al., 2020) dataset, which is a 4-label emotion classification task for a collection of tweets. The prompt template is: <S1> This person feels [MASK]. The label mapping is: anger: angry, joy: happy, optimism: optimistic, sadness: sad. The Huggingface link is: https://huggingface.co/datasets/cardiffnlp/tweet_eval.
TweetEval Hate
(Basile et al., 2019) is a subset of TweetEval (Barbieri et al., 2020) dataset, which is a binary hate speech detection task for a collection of tweets. The prompt template is: <S1> The sentence is [MASK]. The label mapping is: non-hate: neutral, hate: aggressive. The Huggingface link is: https://huggingface.co/datasets/cardiffnlp/tweet_eval.
TweetEval Offensive
(Zampieri et al., 2019) is a subset of TweetEval (Barbieri et al., 2020) dataset, which is a binary offensive text detection task for a collection of tweets. The prompt template is: <S1> It is [MASK]. The label mapping is: non-offensive: polite, offensive: offensive. The Huggingface link is: https://huggingface.co/datasets/cardiffnlp/tweet_eval.
TweetEval Sentiment
(Rosenthal et al., 2017) is a subset of TweetEval (Barbieri et al., 2020) dataset, which is a sentiment analysis task for a collection of tweets. The prompt template is: <S1> This is [MASK]. The label mapping is: negative: terrible, neutral: okay, positive: great. The Huggingface link is: https://huggingface.co/datasets/cardiffnlp/tweet_eval.
TweetEval Irony
(Van Hee et al., 2018) is a subset of TweetEval (Barbieri et al., 2020) dataset, which is a binary ironic text detection task for a collection of tweets. The prompt template is: <S1> The sentence is [MASK]. The label mapping is: non-irony: genuine, irony: sarcastic. The Huggingface link is: https://huggingface.co/datasets/cardiffnlp/tweet_eval.
Rotten Tomatoes
(Pang & Lee, 2005) is a movie review dataset that contains the binary sentiment label of rotten tomatoes movie reviews. The prompt template is: <S1> This is [MASK]. The label mapping is: negative: terrible, positive: great. The Huggingface link is: https://huggingface.co/datasets/cornell-movie-review-data/rotten_tomatoes.
DBpedia14
(Zhang et al., 2015) is a topic classification dataset that contains 14 labels. The prompt template is: <S1> This is about [MASK]. The label mapping is: Company: company, EducationalInstitution: school, Artist: artist, Athlete: sports, OfficeHolder: politics, MeanOfTransportation: transportation, Building: building, NaturalPlace: nature, Village: town, Animal: animal, Plant: plant, Album: music, Film: movie, WrittenWork: book. The Huggingface link is: https://huggingface.co/datasets/fancyzhx/dbpedia_14.
Emotion
(Saravia et al., 2018) is a sentiment analysis dataset with 6 types of emotions of tweets. The prompt template is: <S1> This person feels [MASK]. The label mapping is: joy: happy, sadness: sad, anger: anger, fear: scared, love: love, suprised: shock. The Huggingface link is: https://huggingface.co/datasets/dair-ai/emotion.
20Newsgroups
(Lang, 1995) is a topic classification dataset that contains 20 labels for news type. The prompt template is: <S1> This is about [MASK] news. The label mapping is: alt.atheism: atheism, comp.graphics: graphics, comp.os.ms-windows.misc: windows, comp.sys.ibm.pc.hardware: ibm, comp.sys.mac.hardware: mac, comp.windows.x: windowsX, rec.autos: car, rec.motorcycles: motorcycle, rec.sport.baseball: baseball, rec.sport.hockey: hockey, sci.crypt: cryptography, sci.electronics: electronics, sci.med: health, sci.space: space, misc.forsale: purchase, talk.politics.misc: politics, talk.politics.guns: gun, talk.politics.mideast: mideast, talk.religion.misc: religion, soc.religion.christian: christian. The Huggingface link is: https://huggingface.co/datasets/SetFit/20_newsgroups.
Folio
(Han et al., 2022) is a logic reasoning benchmark with first order logic annotations, and can be reformulated into the 2-sentence natural language inference format. The prompt template is: <S1> [MASK], <S2>. The label mapping is: False: No, Uncertain: Maybe, True: Yes. The Huggingface link is: https://huggingface.co/datasets/tasksource/folio.
Doc NLI
(Yin et al., 2021) is a document level natural language inference task. The prompt template is: <S1> [MASK], <S2>. The label mapping is: not_entailment: No, entailment: Yes. The Huggingface link is: https://huggingface.co/datasets/tasksource/doc-nli.
WANLI
Worker-AI Collaboration for NLI (Liu et al., 2022) is a natural language inference task. The prompt template is: <S1> [MASK], <S2>. The label mapping is: not_entailment: No, neutral: Maybe, entailment: Yes. The Huggingface link is: https://huggingface.co/datasets/alisawuffles/WANLI.
VitaminC
(Schuster et al., 2021) is a task to determine if given evidence supports claims. The prompt template is: <S1> [MASK], <S2>. The label mapping is: REFUTES: No, NOT_ENOUGH_INFO: Maybe, SUPPORTS: Yes. The Huggingface link is: https://huggingface.co/datasets/tals/vitaminc.
bAbI
(Weston et al., 2015) is a set of 20 NLP toy tasks that can be reformulated into natural language inference format. The prompt template is: <S1> [MASK], <S2>. The label mapping is: not_entailment: No, entailment: Yes. The Huggingface link is: https://huggingface.co/datasets/tasksource/babi_nli.
Fake News
is a fake news detection dataset that contains binary labels. The prompt template is <S1> It was [MASK] news. The label mapping is: 0: fake, 1: real. The Huggingface link is: https://huggingface.co/datasets/GonzaloA/fake_news.
Human-vs-Machine
(Nicolai Thorer Sivesind, 2023) is a classification problem to differentiate human and machine generated text. The prompt template is <S1> It was written by [MASK]. The label mapping is: human-produced: human, machine-generated: machine. The Huggingface link is: https://huggingface.co/datasets/NicolaiSivesind/human-vs-machine.
AI-Human-Text
is a classification problem to differentiate human and machine generated text. The prompt template is <S1> It was written by [MASK]. The label mapping is: human-produced: human, machine-generated: machine. The Huggingface link is: https://huggingface.co/datasets/andythetechnerd03/AI-human-text.
WoS
Web of Science (Kowsari et al., 2017) is a topic classification problem that contains 7 labels for paper abstract. The prompt template is: <S1> This is about [MASK]. The label mapping is: Computer Science: CS, Electrical_Engineering: ECE, Psychology: Psychology, Mechanical_Engineering: MechE, Civil_Engineering: CivilE, Material_Engineering: MaterialE. The Huggingface link is: https://huggingface.co/datasets/river-martin/web-of-science-with-label-texts.
PiC
Phrase Similarity (Pham et al., 2022) is the binary classification task to determine whether two phrases are semantically equivalent given the context. The prompt template is: <S1> <S2> Does <PHRASE1> and <PHRASE2> have the same meaning? [MASK]. The label mapping is: negative: No, positive: Yes. The Huggingface link is: https://huggingface.co/datasets/PiC/phrase_similarity.
ART
(Bhagavatula et al., 2020) is the multiple choice task of given observations, determine the most plausible hypothesis from the two. The prompt template is: <S1> A: <Choice1> B: <Choice2> Question: which hypothesis is correct? Answer: [MASK]. The label mapping is: 1: A, 2: B. <S1> is the concatenation of two observations, and choices are two hypotheses. The Huggingface link is: https://huggingface.co/datasets/allenai/art.
ARC
(Clark et al., 2018) is the multiple choice question designed for advanced artificial intelligence reasoning, with Easy and Challenge splits. The prompt template is: <Question> A: <Choice1> B: <Choice2> C: <Choice3> D: <Choice4> E: <Choice5> Answer: [MASK]. The label mapping is: 0: A, 1: B, 2: C, 3: D, 4: E. The Huggingface link is: https://huggingface.co/datasets/allenai/ai2_arc.
HellaSwag
(Zellers et al., 2019) is a multiple choice question task to deterimine which is the best ending of the sentence. The prompt template is: <S1> A: <Choice1> B: <Choice2> C: <Choice3> D: <Choice4> Answer: [MASK]. The label mapping is: 0: A, 1: B, 2: C, 3: D. The Huggingface link is: https://huggingface.co/datasets/Rowan/hellaswag.
PiQA
Physical Interaction: Question Answering (Bisk et al., 2020) is a multiple choice question task that evaluates physical commonsense knowledge. The prompt template is: <S1> A: <Choice1> B: <Choice2> Answer: [MASK]. The label mapping is: 0: A, 1: B. The Huggingface link is: https://huggingface.co/datasets/ybisk/piqa.
SWAG
Situations With Adversarial Generations (Zellers et al., 2018) is a multiple choice question task that evaluates both skill of natural language inference and physical knowledge reasoning. The prompt template is: <S1> A: <Choice1> B: <Choice2> C: <Choice3> D: <Choice4> Answer: [MASK]. The label mapping is: 0: A, 1: B, 2: C, 3: D. The Huggingface link is: https://huggingface.co/datasets/allenai/swag.
SiQA
Social Interaction QA (Sap et al., 2019) is a multiple choice question task that evaluates social common sense knowledge. The prompt template is: <Context> Question: <Question> A: <Choice1> B: <Choice2> C: <Choice3> D: <Choice4> Answer: [MASK]. The label mapping is: 0: A, 1: B, 2: C, 3: D. The Huggingface link is: https://huggingface.co/datasets/allenai/social_i_qa.